Problem 6-3A (Part Level Submission) Windsor, Inc. had a beginning inventory on January 1 of 120 units of Product 4-18-15 at(61) Your answer is correct. Calculate average cost per unit. (Round answer to 2 decimal places, e.g. 2.23.) Average cost per(b2) Determine (1) the ending inventory, and (2) the cost of goods sold under each of the assumed cost flow methods (FIFO, LIhelp me find FIFO and LIFO and AVERAGE COST in part B2
Answer #1
Fifo lifo average cost
The ending inventory

= (280-160)*28 + (80*31)

= 3360+2480

= 5840

= (320-240)*25+(120*22)

= 2000+2640

= 4640

= 200*26.16

= 5232

Cost of Goods Sold

= (120*22)+(320*25)+(200*26)+(160*28)

= 2640+8000+5200+4480

= 20320

= (80*31)+(280*28)+(200*26)+(240*25)

= 2480+7840+5200+6000

= 21,520

= 800*26.16

= 20,928

> where did you get 160 from???

One & Only Fri, Feb 18, 2022 9:17 AM